# Poisoning with A Pill: Circumventing Detection in Federated Learning

### v1

@@ Abstract
Without direct access to the client’s data, federated learning (FL) is well-known for its unique strength in data privacy protection among existing distributed machine learning techniques. However, its distributive and iterative nature makes FL inherently vulnerable to various poisoning attacks, including model poisoning attacks and data poisoning attacks. To counteract these threats, extensive defenses have been proposed to filter out malicious clients, using various detection metrics. Based on our analysis of existing attacks and defenses, we find that there is a lack of attention to model redundancy. In neural networks, various model parameters contribute differently to the model’s performance. However, existing attacks in FL manipulate all the model update parameters with the same strategy, making them easily detectable by common defenses. Meanwhile, the defenses also tend to analyze the overall statistical features of the entire model updates, leaving room for sophisticated attacks. Based on these observations, this paper proposes a generic and attack-agnostic augmentation approach designed to enhance the effectiveness and stealthiness of existing FL poisoning attacks against detection in FL, pointing out the inherent flaws of existing defenses and exposing the necessity of fine-grained FL security. Specifically, we employ a three-stage methodology that strategically constructs, generates, and injects poison (generated by existing attacks) into a pill (a tiny subnet with a novel structure) during the FL training, named as pill construction, pill poisoning, and pill injection accordingly. Extensive experimental results show that FL poisoning attacks enhanced by our method can bypass all the popular defenses, and can gain an up to 7x error rate increase, as well as on average a more than 2x error rate increase on both IID and non-IID data, in both cross-silo and cross-device FL systems.
@@ 1 Introduction
With the prosperity of machine learning and cloud computing, FL [1, 2] has been widely recognized as an effective approach to training machine learning models via distributed data on a large number of scattered clients. Compared with traditional centralized machine learning that collects data to a central cluster, FL does not require direct access to clients’ data, saving the communication overhead while protecting data privacy. However, due to its distributed architecture, FL is inherently vulnerable when clients are compromised by attackers. Numerous studies [3, 4, 5, 6, 7, 8] have demonstrated techniques that manipulate the behavior of trained global model via malicious clients, commonly referred to as poisoning attacks. These attacks can be further divided into the following two types: 1) Model poisoning attacks involve direct modifications of local model updates by the attacker [4, 6]. The corrupted updates, when aggregated, significantly skew the global model’s parameters to the opposite direction. 2) Data poisoning attacks inject malicious samples into the local training datasets [8, 9, 10, 11, 12, 13, 14, 15]. The compromised data samples lead to corrupted local model updates that, once merged, secretly deteriorate the global model’s integrity. Poisoning attacks expose severe security risks to FL [16, 17, 18, 19], violating its integrity and reliability.
To alleviate and further eliminate the impacts caused by poisoning attacks, researchers have proposed various defense methods. They can be divided into four main categories, including adaptive client filtering, statistical parameter aggregation, client-dominant detection, and other advanced metrics and pipelines. Adaptive client filtering [20, 21, 22, 23, 24, 25] utilizes various metrics or scores to evaluate the risk of clients, and discards updates from those with a high risk level. Statistical parameter aggregation [26, 27, 28, 29] modifies the standard aggregation process of FL, granting different weights to different clients based on their model update’s statistical features. Client-dominant detection [30, 31, 32, 33, 34] as a rising category leverages clients to maintain the integrity of FL training. While other advanced metrics and pipelines [35, 36, 37, 38, 39] use other advanced features or pipelines to perform the detection. The effectiveness of existing defense methods hinges on their ability to detect abnormal parameter updates, which are obvious when using most existing attacks. In particular, these attacks treat all neural network parameters equally and manipulate the entire model update using the same mutation strategy.
We argue that it is inefficient and ineffective for attackers to modify all the parameters simultaneously. Researches on model pruning [40, 41, 42, 43, 44] show that model parameters do not contribute equally to a model’s performance. Parameters that have little impact on a model’s performance are considered redundant. Manipulating these redundant parameters not only wastes the attacker’s budget but also decreases the stealthiness of the attack, especially when the FL server analyzes the overall statistics of model updates for anomaly detection. Thus, a more effective strategy is to manipulate critical parameters [45], which are crucial for the model’s performance. This approach ensures that the changes have a significant impact, enhancing the attack’s effectiveness while maintaining stealthiness.
We propose a novel attack-agnostic augmentation method in this paper, enhancing existing model poisoning attacks via a three-stage pipeline, including pill construction, pill poisoning, and pill injection. In the first stage – pill construction, we carefully design an abstract architecture of pill (a tiny subnet with a unique structure), named the pill blueprint, and identify the corresponding pill subnet instance in the target model via a dynamic search algorithm. Then in the second stage – pill poisoning, we reuse existing FL poisoning attacks in an attack-agnostic fashion and concentrate the malicious model updates to the selected pill subnet. Finally, in the pill injection stage, we insert the poisoned pill into a benign model update and use a two-step adjustment to further reduce the difference between the poisoned model update and the original benign update. To this end, we augment existing FL poisoning attacks by dynamically generating a pill, poisoning it, and gradually inserting it into the global model.
@@
We conduct extensive experiments to evaluate the effectiveness of our augmentation method. We use our method to enhance four baseline poisoning attacks, namely, sign-flip attack, trim attack [4], krum attack [4], and min-max attack [6]. Using both the original and augmented attacks, we test the error rates (i.e., the proportion of incorrect predictions) of the global model trained by nine aggregation rules, including FedAvg [2], FLTrust [21], Multi-Krum [20], Median [26], Trim [26], Bulyan [27], FLDetector [39], DnC [6], and Flame [23]. The set of aggregation rules covers most existing defense metrics, to the best of our knowledge. We also design a corresponding adaptive defense to further test the robustness of our method when the defender has the full knowledge of our pipeline and detailed implementation. The experimental results illustrate that our method can substantially improve existing FL poisoning attacks by inducing a more than 2x increase on average in the model prediction error rate under existing defenses, and at most, a more than 7x error rate increase.
Our contributions can be summarized as follows:
@@
  • We propose a generic and attack-agnostic augmentation approach designed to enhance poisoning attacks against robust FL. We are the first to encapsulate model poisoning attacks to well-defined subnets (i.e., pills) combined with comprehensive metric-based performance-boosting adjustment.
  • Extensive results against nine aggregation rules on three common datasets show the attack enhancement capability of our method. In particular, it helps four baseline model poisoning attacks bypass almost all the prevailing defense methods.
@@
  • We observe and point out the inherent limitations of existing model poisoning attacks and defenses, exposing the necessity and potential of fine-grained FL security.
@@ 2.1 Federated Learning
Federated Learning (FL) [1, 2] trains a global model using the information from a swarm of clients without the direct access to each client’s data. In a standard FL training process, within an arbitrary communication round $t$, the FL server first distributes its global model $\boldsymbol{g}_{t}$ to all the clients $K$. After receiving this global model, each client $i$ trains a local model $\boldsymbol{g}_{t}^{(i)}$ with its local data $D^{(i)}$, and uploads the model update $\Delta\boldsymbol{g}_{t}^{(i)}$ to the FL server. After receiving the model updates from the clients, the FL server uses aggregation rules to calculate the global model $\boldsymbol{g}_{t+1}$ for the next round. The objective of FL can be formulated as:
@@ 2.2 Poisoning Attacks in FL
Based on prior investigations [46, 47, 48], existing poisoning attacks in Federated Learning (FL) can be classified according to the techniques employed by attackers. Attackers may directly compromise the global model by manipulating the updates from local models [3, 4, 6, 7, 5] by compromised clients. Alternatively, they may poison their local datasets to indirectly influence the global model [9, 8, 10, 11, 12, 49]. The former technique is referred to as model poisoning attacks in FL, while the latter data poisoning attacks. Although model poisoning attacks are effective, existing attacks have limited stealthiness and can be detected by many existing defenses. Our goal is hence to demonstrate that such attacks can be augmented in a uniform way. Model poisoning attacks directly manipulate the parameters uploaded by clients, with a minimal interference to the local training process. Among these attacks, the simplest form is the sign-flipping attack, which directly flips the model update and scales it with a constant factor. A-Little-is-Enough [3] generates malicious updates within a calculated perturbation range to deceive the global model. Adaptive attacks [4], such as the Trim attack and the Krum attack, dynamically scale malicious updates based on parameter values and distances. The Min-Max and Min-Sum Attacks [6] provide a dynamic scaling for malicious updates based on different distance-based criteria. MPAF [7] aims to drive the global model towards a predefined target model with poor performance on given FL tasks. For the sake of generality, we employ the sign-flipping attack, two types of adaptive attacks [4] (the Trim attack and the Krum attack), and the Min-Max attack [6] as the baseline attacks in this paper.
Additionally, our method’s pill design is inspired by a specialized data poisoning attack known as the subnet replacement attack (SRA) [15]. This approach concentrates backdoor attacks within an arrow-width subnetwork of the original model. It trains this selected subnet using poisoned data and replaces the corresponding parameters of the target model with those from the trained subnetwork. Once the replacement is complete, SRA severs the connections between the poisoned subnetwork and the original model to preserve the efficacy of the attack. The stealthy yet effective design of SRA inspire our method. In particular, we devise a new subnet structure, referred to as the pill blueprint, which features heterogeneous widths to better accommodate a variety of existing FL poisoning attacks. Besides, unlike SRA’s one-time injection, our method gradually poisons the global model throughout the entire FL training, achieving better effectiveness against a wide range of defenses in the FL setting.
@@
2.3 Defenses against Poisoning Attacks in FL
Existing defenses can be categorized based on the mitigation strategies that they utilize.
Adaptive Client Filtering. These techniques such as Krum and Multi-Krum [20] filter out malicious clients through single or multiple rounds of client selection based on distance scores. FLTrust [21] computes trust scores using the cosine similarity between each client update and the server model’s update for weighted averaging. SignGuard [22] employs sign-based clustering combined with norm-based thresholding to identify and filter malicious clients. Flame [23] and Deepsight [24] propose adaptive clustering and clipping to safeguard against backdoor attacks. SkyMask [25] clusters trainable feature masks of clients to assess each client’s risk level.
Statistical Parameter Aggregation. Approaches like Median and Trim [26] use coordinate-wise median or trimmed mean values to aggregate model updates. Bulyan [27] enhances robustness by integrating Krum with Trim techniques. Fool’s Gold [28] applies an adaptive learning rate based on inter-client contribution similarity to mitigate the effects of malicious updates. SparseFed [29] aggregates sparsified updates, reducing the risk of model poisoning attacks.
Client-dominant Detection. Siren [30] and Siren+ [31] set proactive accuracy-based alarms at the client level with the corresponding server-side decisions to counter various model poisoning attacks. FL-WBC [32] introduces client-side noise to diminish the efficacy of attacks and shorten their duration. FLIP [33] achieves higher robustness through client-side reverse-engineering defenses against extensive poisoning strategies. LeadFL [34] uses a client-side Hessian matrix optimization to reduce the impact of adversarial patterns on backdoor and targeted attacks.
Other Advanced Metrics and Pipelines. Various studies employ other sophisticated metric pipelines designed for detection to ensure robust defense against poisoning attacks. These include techniques proposed in studies such as Zeno [35], CRFL [36], FedRecover [37], FLCert [38], FLDetector [39], and MESAS [50].
To comprhensively evaluate our method, we use Multi-Krum (MKrum) [20], Trimmed Mean (Trim) [26], Coordinate-wise Median (Median) [26], Bulyan [27], FLTrust [21], FLDetector (FLD) [39], DnC [6], and Flame [23], a set of representative methods, as our baselines. More details are shown in Appendix A.
3 Threat Model
Attacker’s Goal and Capabilities. This paper focuses on improving the effectiveness of existing poisoning attacks in FL. Similar to previous work [4, 6], an attacker aims to raise the error rates of the global model on a specific class or multiple classes by sending poisoned model updates via compromised clients during the iterative aggregation. Our method does not require any additional knowledge compared with existing FL poisoning attack. Hence we reuse the typical threat model in existing studies [4, 6]. The attacker has a complete control of the compromised clients, including their local data, local training, and uploading process. With the aggregated resources on the compromised clients, the attacker may aggregate the local data from the compromised clients to do extra training or aggregate their local updates to do model estimation. The attacker may or may not know the updates of other benign clients, depending on the confidentiality of the communication channels between the server and clients. Besides, the attacker cannot access the server’s information, including the aggregation rules or the selected clients in each round.
Defense Settings. Most of the defenses in FL are deployed and executed on the server. We adopt a similar defense setting as existing studies [20, 26, 21, 30]. The server cannot directly analyze the local data or the local training of clients. It can only detect malicious clients through model updates from different clients. The server can collect and possess a root test dataset to provide more accurate and robust detection, while the data of such a root test dataset cannot be derived from clients. The data distribution of this root test dataset may or may not be the same as the data distribution across the clients.
@@ 4 Design Objectives and Challenges
  • It presents a significant challenge that the attack augmentation method must use significantly fewer parameters while still achieving similar results as the original attacks.

### v2

@@ Abstract
Federated learning (FL) protects data privacy by enabling distributed model training without direct access to client data. However, its distributed nature makes it vulnerable to model and data poisoning attacks. While numerous defenses filter malicious clients using statistical metrics, they overlook the role of model redundancy, where not all parameters contribute equally to the model and attack performance. Current attacks manipulate all model parameters uniformly, making them more detectable, while defenses focus on the overall statistics of client updates, leaving gaps for more sophisticated attacks. We propose an attack-agnostic augmentation method to enhance the stealthiness and effectiveness of existing poisoning attacks in FL, exposing flaws in current defenses and highlighting the need for fine-grained FL security. Our three-stage methodology, including pill construction, pill poisoning, and pill injection, injects poison into a compact subnet (i.e., pill) of the global model during the iterative FL training. Experimental results show that FL poisoning attacks enhanced by our method can bypass 8 state-of-the-art (SOTA) defenses, gaining an up to 7x error rate increase, as well as on average a more than 2x error rate increase on both IID and non-IID data, in both cross-silo and cross-device FL systems.
Codehttps://github.com/MarkGHX/PoisonPill
@@ 1 Introduction
With the rising demand for machine learning and cloud computing, Federated Learning (FL)(konevcny2016federated; mcmahan2017communication) has emerged as a key approach for training models on distributed data from scattered clients. Unlike centralized machine learning, FL avoids direct data access, reducing communication overhead and enhancing privacy. However, its distributed nature leaves it vulnerable when clients are compromised. Numerous studies (baruch2019little; fang2020local; bhagoji2019analyzing; shejwalkar2021manipulating; cao2022mpaf; bagdasaryan2020backdoor) have examined poisoning attacks, where malicious clients manipulate the global model. These fall into two categories: 1) Model poisoning which directly alters local updates to skew global parameters (fang2020local; shejwalkar2021manipulating); and 2) Data poisoning that injects malicious samples into local datasets (bagdasaryan2020backdoor; tolpegin2020data; xie2020dba; sun2019can; wang2020attack; chen2017targeted; liu2018trojaning; qi2022sra). Such attacks threaten FL’s integrity and reliability (lyu2020threats; kairouz2021advances).
To mitigate these attacks, defenses have been proposed, including adaptive client filtering (krum; cao2020fltrust; xu2021signguard; nguyen2022flame; yan2023skymask), statistical parameter aggregation (yin2018trim; bulyan; foolsgold; panda2022sparsefed; han2023towards), client-dominant detection (guo2021siren; guo2024siren+; sun2021flwbc; zhang2023flip; zhu2023leadfl; park2023feddefender), and other advanced metrics and pipelines (xie2019zeno; xie2021crfl; cao2023fedrecover; cao2022flcert; zhang2022fldetector). These approaches aim to identify suspicious updates, which are usually evident in model poisoning attacks that uniformly alter parameters.
We argue that modifying all parameters uniformly is not a cost-effective approach. Studies on model pruning (frankle2018lottery; lin2018deep; han2015deep; mugunthan2022fedltn; jiang2022model) show that parameters do not contribute equally to a model’s performance. Altering redundant parameters wastes resources and reduces attack stealthiness. A more effective strategy is to target critical parameters (zhang2023oblivion), which significantly impact performance, thereby increasing the attack’s effectiveness while maintaining stealthiness. Thus, we propose a novel attack-agnostic augmentation method that enhances model poisoning attacks using a three-stage pipeline: pill construction, pill poisoning, and pill injection. In the first stage, we design a pill blueprint and identify its corresponding subnet instance in the target model. During pill poisoning, existing FL attacks are applied in an attack-agnostic manner to poison the selected pill. Finally, in pill injection, the poisoned pill is inserted into an estimated benign update, and a two-step adjustment is used to minimize the difference between the poisoned and benign updates. This approach dynamically generates, poisons, and injects a pill into the global model, augmenting existing FL poisoning attacks.
@@
We conduct extensive experiments to evaluate the effectiveness of our augmentation method. We apply it to four baseline poisoning attacks: sign-flipping attack, trim attack (fang2020local), krum attack (fang2020local), and min-max attack (shejwalkar2021manipulating). Using both the original and augmented versions, we measure error rates (i.e., the proportion of incorrect predictions) of the global model trained with nine aggregation rules: FedAvg (mcmahan2017communication), FLTrust (cao2020fltrust), Multi-Krum (krum), Median (yin2018trim), Trim (yin2018trim), Bulyan (bulyan), FLDetector (zhang2022fldetector), DnC (shejwalkar2021manipulating), and Flame (nguyen2022flame). These aggregation rules represent most existing defense metrics. We also design an adaptive defense where the defender has full knowledge of our pipeline and implementation. Results show our method substantially improves existing FL poisoning attacks, leading to over a 2x average increase in model prediction error rates under existing defenses, and up to a 7x increase in some cases.
Our contributions are summarized as follows:
@@
  • We propose a generic, attack-agnostic augmentation method that enhances poisoning attacks against robust FL by encapsulating model poisoning attacks into well-defined subnets (i.e., pills) with comprehensive metric-based adjustments.
  • Extensive experiments on three common datasets against nine aggregation rules demonstrate that our method helps baseline attacks bypass almost all existing defenses, which cannot be attacked by original versions.
@@
  • We identify limitations of existing poisoning attacks and defenses in FL, highlighting the need and potential for fine-grained FL security.
@@ 2.1 Federated Learning
Federated Learning (FL) (konevcny2016federated; mcmahan2017communication) trains a global model using the information from a swarm of clients without the direct access to each client’s data. In a standard FL training process, within an arbitrary communication round $t$, the FL server first distributes its global model $\boldsymbol{g}_{t}$ to all the clients $K$. After receiving this global model, each client $i$ trains a local model $\boldsymbol{g}_{t}^{(i)}$ with its local data $D^{(i)}$, and uploads the model update $\Delta\boldsymbol{g}_{t}^{(i)}$ to the FL server. After receiving the model updates from the clients, the FL server uses aggregation rules to calculate the global model $\boldsymbol{g}_{t+1}$ for the next round. The objective of FL can be formulated as:
@@ 2.2 Poisoning Attacks in FL
Following prior studies (shejwalkar2022back; khan2023pitfalls; jere2020taxonomy), poisoning attacks in Federated Learning (FL) fall into two categories: model poisoning and data poisoning. In model poisoning attacks, adversaries compromise the global model by directly altering local model updates (baruch2019little; fang2020local; shejwalkar2021manipulating; cao2022mpaf; bhagoji2019analyzing). In data poisoning attacks, they corrupt local datasets to indirectly affect the global model (tolpegin2020data; bagdasaryan2020backdoor; xie2020dba; sun2019can; wang2020attack; zhang2022neurotoxin). Further details are in the extended version.
Our pill design draws inspiration from the subnet replacement attack (SRA) (qi2022sra), a backdoor injection method that limits backdoors to a small subnetwork. SRA trains this subnet with tainted data, substitutes the target model’s parameters, and disconnects the subnet to maintain attack effect. Drawing from SRA’s stealthy yet potent design, we propose a heterogeneous-width pill blueprint for varied FL poisoning attacks. Unlike SRA’s one-time injection, our approach gradually poisons the global model during training, improving robustness against various defenses.
2.3 Defenses against Poisoning Attacks in FL
Existing defenses can be categorized based on the mitigation strategies that they utilize, including Adaptive Client Filtering, Statistical Parameter Aggregation, Client-dominant Detection, and Other Advanced Metrics and Pipelines. To comprhensively evaluate our method, we use Multi-Krum (MKrum) (krum), Trimmed Mean (Trim) (yin2018trim), Coordinate-wise Median (Median) (yin2018trim), Bulyan (bulyan), FLTrust (cao2020fltrust), FLDetector (FLD) (zhang2022fldetector), DnC (shejwalkar2021manipulating), and Flame (nguyen2022flame), a set of representative defenses, as our baselines. More details are in the extended version.
2.4 Threat Model
We follow the typical threat model used in existing studies (fang2020local; shejwalkar2021manipulating), where the attacker has access to a subset of compromised clients and aims to increase the error rates of the global model on specific classes or across all classes. In this scenario, defenses cannot directly analyze the data on each client as the defender’s setting in  (krum; yin2018trim; cao2020fltrust; guo2021siren). Instead, they identify malicious clients by analyzing the uploaded client updates. Further details are in the extended version.
@@ 3 Design Objectives and Challenges
  • It presents a significant challenge that the attack augmentation method must use significantly fewer parameters while still achieving similar attack effects.